Translate the sentence below and write the 2 equations. Then solve the equations using substitution or addition.
Six bananas and three kiwis cost $3.36. Two bananas and five kiwis cost $3.04. Find the cost of a banana and the cost of a kiwi.
b=
k=

Found 2 solutions by stanbon, unlockmath:
Answer by stanbon(48558)

Translate the sentence below and write the 2 equations. Then solve the equations using substitution or addition.
----------------------------------------------------
Six bananas and three kiwis cost $3.36. Two bananas and five kiwis cost $3.04. Find the cost of a banana and the cost of a kiwi.
-------------
Equations:
6b + 3k = 336 cents
2b + 5k = 304 cents
-----------------------------
Multiply 2nd equation by 3 to get:
6b + 15k = 912
---
Subtract that eq. from the 1st and solve for "k":
-12k = -576
k = 48 cents (price of a kiwi)
----------------------
Substitute that value into 2b + 5k = 304 to solve for "b":
2b + 5*48 = 304
2b = 64
b = 32 cents (price of a banana)
===================================
Cheers,
Stan H.

Answer by unlockmath(1122)

Hello,
Set up the following.
6 times b would give the cost of the bananas where b is the price of one banana.
3 times k would be the cost for the kiwis. Total is $3.36 The equation is:
6b + 3k = $3.36
The other equation would look like this:
2b + 5k = $3.04 Let's multiply this equation by -3 so to get 2b into -6b then we'll add both equations together.
6b + 3k = $3.36
-6b -15k = -9.12
We add the two above equations to get:
-12k = -5.76 Solve for k by dividing -12 into both sides.
k = .48 So the kiwi costs $0.48 each. Plug this into one of the original equations to look like this:
6b + 3(.48) = $3.36 Subtract $1.44 from both sides.
6b = $1.92 Divide each side by 6
b = $.32 So the bananas each cost $0.32
Kiwis cost $0.48 each
Bananas cost $0.32 each
RJ Toftness
www.math-unlock.com